OpenStudy (anonymous):

MEDAL given work two problems out pleaseee 1. x^2- 8x -7=0 2. x^2 -3x=40

OpenStudy (campbell_st):

well if you are solving them for question 1, you'll need the general formula \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] and a = 1, b = -8 and c = -7 just substitute and evaluate Question 2. rewrite the equation \[x^2 - 3x - 40 = 0\] this can be solved by factoring what are the factors of -40 that add to -3.... the larger factor is negative then the factored form is (x + factor 1)(x + factor 2) = 0 then you need to solve x + factor 1 = 0 x + factor 2 = 0 hope it helps.

OpenStudy (anonymous):

@surjithayer can you help with two problems pleaseeee

OpenStudy (anonymous):

1. x^2-8x-7=0 x=[-(-8)+-sqrt{(-8)^2-4*1*-7}]/2*1 ={8+-sqrt92}/2 =(8+-2sqrt23)/2 =4+-sqrt23 Hence x=4+sqrt23 and x=4-sqrt23

OpenStudy (campbell_st):

lol... why give answers... how does it help understanding...

OpenStudy (anonymous):

2. x^2-3x-40=0 1*-40=-40 8-5=3 8*-5=-40 write -3x=-(8-5)x then make factors.

OpenStudy (anonymous):

@surjithayer you did the good like always thanks!

OpenStudy (anonymous):

@campbell_st actually i have the answer i just need to work the problem out so whatever!

OpenStudy (campbell_st):

lol... so why not post your answers with the questions... it makes the process of checking so much easier...

OpenStudy (campbell_st):

and a very simple way to check you solutions is to substitute back into your equation... and see if you get the desired answer... zero...

OpenStudy (anonymous):

thats true thankssss @campbell_st